An auto manufacturer sends cars from its two plants, Plant I and Plant II, to two delearships, dealership A and delearship B. Plant I ships 85 cars to the delearships and Plant II ships 70 cars. Dealership A needs 53 cars and dealership B needs 102 cars. The cost for shipping a car from Plant I to dealerships A and B is $310 and $220, respectively. The cost for shipping a car from Plant II to dealerships A and B is $330 and $150, respectively. There are $35050 available for shipping the cars.

An auto manufacturer sends cars from its two plants, Plant I and Plant II, to two delearships, dealership A and delearship B.Let x represent the number of cars shipped from Plant I to dealership A. Let y represent the number of cars shipped from Plant II to dealership A. Let z represent the number of cars shipped from Plant I to dealership B. Let w represent the number of cars shipped from Plant II to dealership B.
Answer #1

Consider the equations

x+y=53    equation 1      

z+w=102 equation 2

x+z=85 equation 3

y+w=70 equation4

310x+220z+330y+150w=35050 equation5

(Equation 4)-(equation 1)

\Rightarrow y+w-(x+y)=70-53

\Rightarrow y+w-x-y=70-53

w-x=17

w=17+x equation a

(equation 3)-(equation1)

\Rightarrow x+z-(x+y)=85-53

\Rightarrow x+z-x-y=85-53

\Rightarrow z-y=32

z=32+y equation b

consider equation 5

\Rightarrow 310x+220z+330y+150w=35050

\Rightarrow 31x+22z+33y+15w=3505

substitute w=17+x and z=32+y

\Rightarrow 31x+22(32+y)+33y+15(17+x)=3505

\Rightarrow 31x+704+22y+33y+255+15x=3505

\Rightarrow 46x+55y=3505-(704+255)

\Rightarrow 46x+55y=2546 equation 6

consider equatioequn 1

x+y=53

multiplying this equation by 46 gives

46x+46y=53*46

46x+46y=2438 equation 7

(equation 6)-(equation 7)

46x+55y-46x-46y=2546-2438

55y-46y=108

9y=108

\Rightarrow y=12

y is the number of cars shipped from plant II to dealership A

\therefore\;12\;cars\;are\;shipped\;from\;plantII\;to\;dealership\;A

Substitute y=12 in equation 1

x+y=53

x+12=53

x=53-12

x=41

x is the number of cars shipped from plant I to dealership A

Therefore 41 cars are shipped from plant I to dealership A

Substitute x=41 in equation 3

x+z=85

41+z=85

z=85-41

z=44

z represent number of cars shipped from plant I to dealership B

Therefore 44 cars are shipped from plant I to dealership B

Substitute z=44 in equation 2

z+w=102

44+w=102

w=102-44

w=58

Number of cars shipped from plant II to dealership B is 58